A manufacturer produces lots of a canned food product. Let p denote the proportion of the lots that do not meet the product quality specifications. An

n = 37, c = 0

acceptance sampling plan will be used.

(a)

Compute points on the operating characteristic curve when

p = 0.01, 0.03, 0.10, and 0.20.

(Round your answers to four decimal places.)

c

p = 0.01

p = 0.03

p = 0.10

p = 0.20

0            

(b)

Plot the operating characteristic curve.


(c) What is the probability that the acceptance sampling plan will reject a lot containing 0.10 defective? (Round your answer to four decimal places.)
Answer #1

Answer: Given n=37 c=0 / using By nex pt qnx FOD P=0.01- PC x < 0) PCX co2 *00-01) (0.99337-0 = 3700 – IX IX 0.6894 0:6894 P=b)

1.00 D.RU 0.60+ Probability of Accepting the lat 0 40+ 0.20 0.00 20 5 10 15 Percent Defective in the Lot

acceptance c) the probabisity that the sampling pian will beject a lot containing 0.10 defective : 1- P(x<o) - (-2CX- =1-(37c